OpenStudy (anonymous):

HELP WANTED!!!! USE THE FACT THAT THE TRIGONOMETRIC FUNCTIONS ARE PERIODIC TO FIND THE EXACT VALUE OF THE GIVEN EXPRESSION. TAN (405° )

OpenStudy (anonymous):

405 degrees is coterminal with 45 degrees because 360+45=405

OpenStudy (anonymous):

I don't know how to use periodicity, but you can use the formula\[\large \tan(a+b)=\frac{tana+tanb}{1-tana\times tanb}\]

OpenStudy (anonymous):

Oh ok that makes sense :)

OpenStudy (anonymous):

so now , tan(405) = tan(45)

OpenStudy (anonymous):

or you can do that ^^^

OpenStudy (anonymous):

absolutely nothing wrong with that...

OpenStudy (anonymous):

Yeah but your solution is way easier and faster..and actually does what the question expects you to do :D

OpenStudy (anonymous):

ALRIGHTY THANKS GUYS

OpenStudy (anonymous):

yw...